### v2

# Fairly Accurate: Fairness-aware Multi-group Target Detection in Online Discussion

Abstract
Target-group detection is the task of detecting which group(s) a social media post is “directed at or about”, with various applications, such as targeted-marketing. In this work, we focus on the fairness implications of target-group detection in the context of toxicity detection, where the perceived harm of a post often depends on which group(s) it targets. Because toxicity is highly contextual, language that appears benign in general may be harmful when targeting specific demographic groups. It is thus important to first detect which group(s) are being targeted by a post as a precursor to the subsequent task of determining whether the post is toxic given the group(s). Target-group detection is also challenging: a single post may simultaneously target one to many groups, and we must detect groups fairly in order to promote equitable treatment. We show that our proposed approach to fairness-aware multi target-group detection not only reduces bias across groups, but also achieves competitive predictive performance, outperforming existing fairness-aware baselines. To spur future research on fairness-aware target-group detection and support competitive benchmarking, we also share our code.
Keywords: Target-group Detection, Accuracy Parity, Group Accuracy Parity, Impossibility Theorem.
1 Introduction
In online discussions, it is important to understand which groups(s) are being referred to in order to properly interpret the discourse. Target-group detection is the task of detecting which group(s) a social media post is “directed at or about” [38], with downstream applications such as targeted-marketing and toxicity detection. In the latter case of toxicity detection, one asks whether a post is hateful or discriminatory toward specific demographic groups (vs. a task such as offensive language detection [12, 49], where offensiveness does not involve specific demographics). Toxicity is also often highly contextual, with language that might seem benign taking on harmful undertones when it targets specific groups [11, 40, 29, 24, 6]. Target-group detection thus represents an important, distinct, and separable task from any downstream tasks based on detected targets.
A key aspect of target-group detection is that a single post may simultaneously target one to many groups. However, prior work has predominantly framed target-group detection as a single-label task, assuming each post targets at most one group [5, 40]. Moreover, most toxicity detection datasets [47, 12] largely lack granular demographic annotations or focus narrowly on binary classifications, failing to account for the complex interplay of cultural, linguistic, and identity-based differences in what constitutes harmful content. Simplifying the task in this way represents a natural starting point for research but misses the multi-label nature of real-world discourse, where harmful language can implicate multiple groups simultaneously.
Another important consideration is that it is not enough to detect groups accurately; we must also detect groups fairly in order to promote equitable treatment across diverse user demographics. For example, in targeted-marketing of appropriate pharmaceutical drugs to patient populations with different diseases, both groups are equally important to detect because each deserves the same opportunity to benefit from the appropriate drug. Detection models trained without fairness constraints may underperform on marginalized groups, reinforcing existing societal biases or harms and leading to disproportionate impact in downstream applications.
Post Toxic Demographic Groups Targeted (i.e., group(s) the post is “directed at or about”)
Asian Black Latinx Middle Eastern Native American Pacific Islander White
I’m chinese but have really dark skin due to my parent’s genes, and I agree, my country’s (singapore) people are all friends with different races and religions.
There is plenty of evidence that women and non-white people are disadvantaged in almost every hiring process in the western world.
… go back to the sh** hole you crawled out of. You America hating jihadi.
Table 1: Sample comments from the MHS Corpus [38], showing the multi-label format for groups. Note that groups can be targeted by a post regardless of the post’s toxicity. To minimize harm and provide a content warning to readers, we have partially redacted toxic language in this table (e.g., profanity masked with asterisks or ellipses).
To date, most fairness literature has focused on decision-making scenarios for allocating goods or burdens, such as college admissions or bail decisions, where false positives and false negatives have asymmetric costs [26]. Target-group detection is notably different. With the pharmaceutical drug example above, misidentifying either group would be equally harmful since either would miss the opportunity to benefit from the appropriate drug for their disease. This motivates symmetric error weighting, promoting equitable treatment of groups and mitigating biases that could amplify disparate impact.
When considering target-group detection for toxicity detection, symmetric error weighing is also desirable. Toxicity detection itself can be viewed as a task of allocating goods and burdens: from the perspective of targets, content moderation that protects the target from toxic content is a good, and therefore gaps in false negatives in toxicity detection may constitute an issue of distributive justice [35, 20]. In contrast, target-group detection is a multi-label task wherein labels are demographic groups themselves (e.g., Latinx, Black), and both error types, misidentifying one group for the other or vice-versa, are equally harmful. For instance, mistaking a post targeting Latinx as one targeting Black is as undesirable as the reverse.
In this work on fairness-aware target-group detection, we address the two key aspects above: that a single post may target one to many groups (e.g., see Table 1); and that detection must be conducted fairly to promote equitable treatment of groups. We use the recent, large-scale Measuring Hate Speech (MHS) Corpus [38] that labels granular information about the type of content in a post (e.g., inciting violence, humiliating, etc.), as well as identifying all groups targeted by each post. Crucially, note that targets are identified for both toxic and non-toxic posts; a social media post can be “directed at or about” [38] one or more target group(s) without being toxic. This dataset lets us to move beyond simplistic definitions of targeted groups by encompassing diverse demographic dimensions: race, gender, religion, and other axes of identity.
In regard to the symmetric cost of group-level errors in target-group detection, Accuracy Parity (AP) [54] provides a natural fairness criterion in seeking balanced predictive accuracy across groups to mitigate bias. To optimize for AP during model training, we adopt and extend the recently proposed Group Accuracy Parity (GAP) loss [19], which penalizes accuracy deviations across groups. Specifically, because the existing GAP formulation is limited to binary settings and does not scale to multi-label scenarios, we introduce $GAP_{multi}$, a scalable loss function that extends GAP to enable fair detection when posts target multiple groups. Our formulation enables fair optimization in realistic multi-group settings while preserving the convergence behavior of the original GAP formulation, thereby ensuring group fairness and global utility.
While Equalized Odds [21] could be considered as an alternative to AP to independently enforce equality in both error types, EO does not actually treat both types of errors equally, and its use in target detection setting could lead to unintended consequences. Through a formal theoretical analysis, we present an impossibility result: under realistic scenarios (e.g., unequal group base rates), one cannot simultaneously satisfy both AP and EO. Moreover, we show that enforcing EO can actually degrade performance for some groups. These findings reinforce AP as a more appropriate fairness objective for our task.
We evaluate multi-group $GAP_{multi}$ loss against a set of baseline loss functions, including traditional fairness-aware losses and standard optimization strategies. We assess two key objectives: balanced performance across demographic groups (group-fairness) and overall predictive accuracy (utility). $GAP_{multi}$ not only achieves better group-fairness (by minimizing performance disparities impact across groups), but also maintains competitive overall accuracy. Our work thus effectively addresses the unique challenges of multi target-group detection, consistently reducing bias across groups while maintaining similar overall utility.
To summarize, we propose a fairness-aware target-group detection framework that optimizes for both utility and group fairness in a multi-label setting. Jointly optimizing for group fairness and utility supports making predictions that are both fair and accurate, addressing challenges posed by imbalanced datasets while capturing group-specific biases. Our work bridges theoretical fairness objectives with real-world applicability, offering a robust and practical solution for group-identification tasks. To spur future research on fairness-aware target-group detection and support competitive benchmarking, we also share our code [^1] URL: https://github.com/smjtgupta/GAP.
2 Related Work
Fairness in algorithmic decision making.
Fairness in machine learning has been extensively studied across many critical applications such as risk assessment [21], college admissions [26], loan approval [9], and recidivism [2]. These domains typically involve tasks where asymmetric errors, i.e., allocating goods or burdens, are central to the decision-making process. For instance, in loan approvals, false negatives (denying credit to a qualified individual) may carry a greater societal cost than false positives (approving a loan for someone who defaults). Similarly, in recidivism prediction, false positives (predicting someone will reoffend when they would not) disproportionately impacts marginalized groups, making error prioritization critical.
For this reason, fairness measures such as Equalized Odds [21] and Demographic Parity [53] are popular. Due to their focus on error rate alignment across groups, these measures are particularly suited for tasks where the interpretation of error rates is tied to the societal implications of unequal access to resources or unjust burdens. In contrast, we investigate multi target-group detection, which presents a fairness challenge where error types are symmetric: misidentifying a targeted group (false negative) and incorrectly assigning a group (false positive) are equally undesirable. This symmetry is relatively rare in traditional fairness applications but critical here, motivating the use of Accuracy Parity (AP) [54] as a natural and principled measure. While AP has been studied conceptually [3, 10, 32], relatively little work has explored its role in multi-group detection settings.
Target-group detection in Toxicity modeling.
Identifying which demographic group(s) are targeted by toxic language has become a crucial component in modern toxicity detection pipelines. Early work by Blodgett et al. [5] framed target-group detection as a single-label classification task across four groups, selecting the most likely group during inference. Similarly, many datasets [12, 16] treat group identification as binary or mutually exclusive, assuming each post targets only one group. However, real-world toxic content often implicates multiple communities simultaneously. While recent studies discuss the need the considering group-targeted toxicity [45] or seek to model it [28, 18], to the best of our knowledge, all have assumed oracle detection of target groups based on annotated corpora.
Fairness-aware loss functions.
A central challenge in fairness-aware machine learning is aligning the training objective (i.e., the loss function) with the fairness criterion used for evaluation. When this alignment is weak, as in the case of many surrogate loss functions [50, 51, 7], models may experience metric divergence, especially under sensitive hyperparameter regimes [33, 31, 52, 44]. These surrogate losses typically optimize for proxies (e.g., cross-entropy with reweighting) rather than directly minimizing fairness violations. In contrast, strictly-mapped loss functions [19, 42] are constructed to directly correspond to fairness measures. These formulations offer a smoother and more reliable optimization path — as the loss decreases, the corresponding fairness metric improves predictably.
Datasets for Target-Group Detection.
The development of fairness-aware target-group detection models has been hindered by the lack of datasets that provide fine-grained, multi-label annotations of demographic groups targeted in social media posts. Early toxicity datasets [47, 12] label content with binary hate labels and limited demographic group references. Subsequent corpora, including those by Blodgett et al. [5], Founta et al. [17] similarly fail to represent the complex overlap of targeted identities present in real-world discourse. Moreover, few datasets enable robust fairness evaluation, as they lack annotator consensus indicators, group-level prevalence statistics, or non-toxic examples that still target identity groups. To address these gaps, we adopt the MHS Corpus [38], a large-scale dataset that explicitly annotates several demographic groups and captures both toxic and non-toxic instances of group references. The dataset supports multi-label annotations per post making it uniquely suitable.
3 Multi Target-group Detection
Group-specific toxicity often manifests through distinct language patterns: coded language may be used against certain ethnicities and innocuous terms can become racialized slurs in certain contexts. Certain forms of toxicity such as gender-based hostility often also involve stereotype-driven insults or harassment. Such content, though superficially benign, may acquire harmful undertones when directed at particular communities. Without explicitly detecting the intended target group(s), toxicity detection models may risk generalizing prediction, leading to misclassifications and disparate impact for the statistical minority groups [40, 11].
Use of annotator-provided target labels enables training of group-conditioned toxicity detection models [18, 15, 20], capturing nuanced differences between shared and group-specific toxicity patterns. However, deploying such systems in practice requires solving the real-world task of target group detection in the wild. Unlike controlled annotation settings, real-world content typically lacks annotated labels identifying target groups; user reports may provide this information, but proactive toxicity detection prompts the need for an accurate and fair target-detection module.
3.1 Problem Statement
To formally define the fairness-aware multi target-group detection task, we assume a dataset $X\in\mathbb{R}^{N\times F}$, where each of the $N$ samples represent a post, and $F$ denotes the dimensionality of extracted features. Let $G$ be the total number of demographic groups. Let $Y\in\mathbb{R}^{N\times G}$ denote the label space, where $y=\{0,1\}^{G}$ represents the set of $G$ binary labels, indicating whether each group $g\in G$ is targeted ($1$) or not (00) by the post $x$. This leads to a multi-label classification problem setting [30, 23], since a post $x$ can target multiple groups simultaneously. The objective is to learn a prediction function set $f:X\rightarrow\{0,1\}^{G}$ that maps each post $x$ to its corresponding multi-group label $y$. For each group $g$, we define the group-specific and overall model accuracy as $Acc_{g}$ and $Acc_{overall}$ respectively.
$Acc_{g}$ $=\frac{\textrm{Number of correct predictions for group }g}{\textrm{Total number of predictions for group }g}$ (1)
$Acc_{overall}$ $=\frac{\sum_{g=1}^{G}\textrm{Number of correct predictions for group }g}{\textrm{Total number of predictions over all groups }G}$ (2)
From an overall utility perspective, we aim to maximize overall accuracy across all groups $G$. However, to ensure fairness, we also seek to minimize disparities in predictive accuracy across groups. We define the inter-group disparity as the deviation of each group’s accuracy from the mean overall accuracy. Thus, we have the following set of optimization problems in Eq. 3 and Eq. 4:
[…]
To jointly optimize for both utility and fairness, we adopt a Single Objective Optimization (SOO) goal, combining these objectives, Eq. 3 as primary and Eq. 4 as secondary, with a regularization term $\lambda\in(0,\infty)$ in Eq. 5. This formulation explicitly balances model utility with group-level fairness, penalizing uneven performance across groups.
[…]
4 From Measure to Optimizer: AP & GAP
Achieving fairness in target-group detection requires not only principled metrics to measure fairness but also optimization strategies that enforce these fairness objectives. In this section, we discuss both Accuracy Parity (AP) [54] as a fairness evaluation measure for target-group detection, as well as Group Accuracy Parity (GAP) [19] as its optimization-aligned counterpart.
Accuracy Parity (AP)
To address the trade-off between accuracy vs. group fairness, AP measures the degree to which predictive accuracy varies across demographic groups. It is suitable for tasks with symmetric error costs, such as target-group detection. AP is well-studied under various names: Accuracy Equity [13], Equality of Accuracy [22], Accuracy Parity [54], Equal Accuracy [32], Overall Accuracy Equality [3] and Accuracy Difference [10].
To date, AP has predominantly been discussed as an evaluation measure, while less effort has been devoted to developing suitable loss functions. This may stem in part from the fact that a bulk of algorithmic fairness literature has been devoted to the use of ML in the allocation of goods and burdens, where this is typically not deemed a desirable fairness measure since it does not capture the different costs of different types of errors. For multi target-group detection, AP is particularly well-suited because it aligns with the symmetric error costs inherent in the task. In this setting (as defined in Eq. 4), errors from all groups are treated equally, irrespective of group membership. For instance, misclassifying a post targeting one group as targeting another group is equally undesirable, regardless of the directionality of the error. A particular attribute of the target-group detection task also merits note: the group and the label are the same. AP thus measures the extent to which all groups are treated equitably.
Group Accuracy Parity (GAP)
To directly optimize for Accuracy Parity (AP), we adopt Group Accuracy Parity (GAP) loss [19], the first differentiable loss function for AP that we are aware of. Unlike surrogate losses that loosely correlate with fairness metrics, GAP strives to minimize the loss directly to improve AP during training. The GAP formulation shown in Eq. 6 combines the overall error (OE) with a regularization term measuring the disparities in cross entropy (CE) error across the binary groups ($g=0,1$). It allows different weighted variants of accuracy w.r.t. chosen entropy, depending on the evaluation need.
$GAP=OE+\lambda\|\underbrace{CE(g=1)}_{\text{err Group 1}}-\underbrace{CE(g=0)}_{\text{err Group 0}}\|_{2}^{2}$ (6)
The regularization term evaluates to zero when the errors across the two groups are same, i.e., perfect AP score of 0. The squared two-norm formulation supports a smooth loss surface that results in better optimization trajectory. While formulated for binary group settings, this structure forms the foundation for our multi-label extension, as discussed next.
5 GAP for Multiple Targets - $GAP_{multi}$
To account for the multi-label nature of target-group detection, we propose a mathematical extension of GAP loss ($GAP_{multi}$) to handle multi-label settings. In our proposed setting, a post can target none, one, or many to all of the groups simultaneously. Our $GAP_{multi}$ loss in Eq. 7 retains the overall error (OE) term (as in Eq. 5) to preserve utility, but it augments OE with a pairwise regularization component that minimizes squared differences in cross-entropy errors between all pairs of demographic groups. This formulation avoids dependence on a global mean error (as used in Eq. 5), enabling parallelizable fairness enforcement across $\prescript{G\mkern-0.5mu}{}{C}_{2}$ group pairs. The result is a scalable and theoretically grounded loss function that directly supports fairness in complex, multi-target classification scenarios.
While we could have followed the Single Objective Optimization format in Eq. 5, there are some implementation issues associated with it. Decomposing it, we can see that the Overall loss in Eq. 3 can be computed in parallel over all the groups $G$. Now one has to wait for all these parallel computations to be over, before coming to Eq. 4, the serial bottleneck component of measuring the mean error value. Only once this mean error is computed can the group-level deviations be measured. This deviation-from-mean computational approach provides a serial component acting as a bottleneck in the otherwise parallel pipeline. To bypass this serial component bottleneck and improve parallelization, we formulate our $GAP_{multi}$ in the form in Eq. 7.
$GAP_{multi}=OE+\lambda\underset{i,j\in[G],i\neq j}{\sum}\|\underbrace{CE(g=i)}_{\text{err Group i}}-\underbrace{CE(g=j)}_{\text{err Group j}}\|_{2}^{2}$ (7)
The regularization term in Eq. 7 considers pairwise error differences $\|CE(g=i)-CE(g=j)\|$ for all distinct $\prescript{G\mkern-0.5mu}{}{C}_{2}$ group pairs $i,j\in[G],i\neq j$. This allows for direct comparisons between specific groups, ensuring that no group is unduly favored or marginalized relative to another. In contrast, deviation-from-mean approaches measure errors against a global aggregate (Eq. 4), which can mask disparities between individual groups. For instance, significant differences in performance between the statistical minority groups may remain undetected if their combined performance aligns with the statistical majority group’s average error. The pairwise structure enforces symmetric treatment of errors across all groups. This is critical for multi-label tasks where multiple groups can be targeted simultaneously.
$GAP_{multi}$ has $\prescript{G\mkern-0.5mu}{}{C}_{2}$ or $G(G-1)/2$ regularization terms. These pairwise error computations, despite having quadratic growth with the group cardinality $G$, are independent and can be performed in parallel. The formulation is thus computationally scalable, even for large datasets with many demographic groups. Global deviation-based approaches rely on mean computation, as serial component, therefore has bottlenecks associated with their sequential nature. In contrast, the $GAP_{multi}$ formulation allows for parallel computations of error terms over arbitrarily large $G$ without any such bottlenecks. Thus it can be integrated seamlessly into modern deep learning pipelines for efficient optimization.
Finally, the pairwise design of $GAP_{multi}$ enables the model to capture fine-grained disparities in predictive performance across specific demographic group pairs. The formulation enforces that error disparities between two targeted groups are directly addressed, enabling a more equitable treatment. This granularity is absent in deviation-from-mean approaches, which treat all groups as equally related to the mean, overlooking important group-specific dynamics of group disparities. This makes our $GAP_{multi}$ formulation well-suited for multi-label scenarios where performance disparities are not evenly distributed, providing both scalability and interpretability for fairness-sensitive applications.
5.1 Code Flow
Alg. 1 presents the logic of the overall weighted Binary Cross Entropy (WBCE) loss, where the weights $w[g]$ represent inversely scaled values of labels (0s and 1s), within each group $g$. Thus $err\_grp[g]$ equivalently maps to Balanced Accuracy (BA), an evaluation measure widely used in datasets with label imbalance. Absence of this weighting term would strictly map to standard accuracy, following the mapping from $BCE$. The overall error $err\_overall$ is then defined as summation of errors across groups. Note that we do not weight groups while adding their errors since: a) we want to treat all groups equally; and b) TensorFlow’s $bce$ function is scale independent, i.e., it produces same error value for equal ratios of mispredictions w.r.t. total samples, irrespective of sample size. For e.g., $bce$ value over 5 samples with 1 misprediction is equal to $bce$ value over 15 samples with 3 misprediction.
Algorithm 1 $overall\_loss(y\_true,y\_pred,w[g])$
Loss function for optimizing Overall Error
1:Input: $y\_true,y\_pred$triangleright True and Predicted Labels
2:Input: $w[g]$triangleright Balanced weights of Group $g$
3:$y\_true\_lab=y\_true[:\,,0]$triangleright Label Info
4:$y\_true\_dem=y\_true[:\,,1]$triangleright Demographic Info
5:for each group $grp[g]\in G$ do
6:     $pos\_grp[g]=group(y\_true\_dem==g)$triangleright Find indices
7:     $y\_true[g]=y\_true\_lab[pos\_grp[g]]$triangleright True group labels
8:     $y\_pred[g]=y\_pred[pos\_grp[g]]$triangleright Predicted group labels
9:     $err\_grp[g]=bce(y\_true[g],y\_pred[g],w[g])$triangleright wBCE
10:$err\_overall=\sum^{G}err\_grp[g]$triangleright summation of group loss
11:Output: $err\_overall$
Alg. 2 presents the logic of our proposed $GAP_{multi}$ loss. After computing overall error via Alg. 1, $GAP_{multi}$ computes the $\prescript{G\mkern-0.5mu}{}{C}_{2}$ group-pairwise errors. The errors are squared to enforce positive values and allow for a smooth loss surface. The final SOO error as per Eq. 5 is the summation of the overall error and a regularized sum of group-pair errors.
Algorithm 2 $gap\_loss(y\_true,y\_pred,w[g])$
Loss function for optimizing Group Balanced Error
0:Repeat Steps 1-10 of Alg. 1
1:for group pairs $[i,j]\in G,[i\neq j]$ do $\triangleright$ $\prescript{G\mkern-0.5mu}{}{C}_{2}$ iterations
2:     $err\_group\_pairs=\sum(err\_grp[i]-err\_grp[j])^{2}$
3:$err\_balanced=err\_overall+\lambda\cdot err\_group\_pairs$
4:Output: $err\_balanced$
6 Impossibility Results
As discussed earlier, achieving fairness in target-group detection requires accounting for both false positives (type-I errors) and false negatives (type-II errors) because both can represent harm to marginalized groups. One thus might consider Equalized Odds (EO) [21] or its surrogates [51, 7] as a viable optimizer. However, as noted in Section 2, EO is not well-suited for this task because it assumes asymmetric error costs. We now present a theoretical impossibility result demonstrating that EO and Accuracy Parity (AP) cannot generally be satisfied simultaneously unless strong assumptions (e.g., equal base rates) hold. This challenges the common misconception that EO implicitly ensures parity in group-level predictive accuracy.
The theorems are stated below, with their corresponding proofs in Appendix A. Furthermore, we highlight that in datasets with base rate disparities, enforcing EO might severely reduce the predictive accuracy for statistical minorities, at the cost of balancing error rates. For nomenclature, let $\{A\}$ be a demographic group having $P_{A}$ positive and $N_{A}$ negative samples, and let the proportion of positive to negative examples in a group be defined as $\delta_{A}^{*}=\frac{P_{A}}{N_{A}}$. For target-group detection, a positive / negative label for each group indicates that group being targeted / not targeted. Thus a false positive (or false alarm) indicates incorrectly detecting a group as being targeted, while a false negative (or miss) indicates failing to detect a group as being targeted.
Theorem 1.
A fairness problem aiming to simultaneously satisfy Equalized Odds and Accuracy Parity is only feasible when: 1. the base rates are equal across all groups; or 2. the model engages in random prediction.
This theorem highlights the limitations in attempting to satisfy both EO and AP simultaneously. When base rates differ across groups, a common scenario in real-world datasets, achieving parity in error rates (EO) inherently conflicts with maintaining balanced predictive accuracy (AP). The only scenarios where these objectives can coexist are either under the idealized condition of equal base rates ($\delta^{*}_{A}=\delta^{*}_{B}$), which rarely holds in practice, or when the model resorts to random predictions, which sacrifices utility altogether.
Theorem 2.
A fairness problem aiming to simultaneously satisfy False Positive and Negative Error Disparity (FPNED) and Accuracy Parity is only feasible when the base rates are equal across groups.
This theorem states that the alternative fairness measure FPNED [7] is incompatible with AP as well for most real-world datasets when base rates are unequal across groups ($\delta^{*}_{A}\neq\delta^{*}_{B}$). The constraint arises because FPNED requires equal error rates (FPR/FNR) across groups, while AP demands equal accuracy—conditions that cannot hold simultaneously under unequal base rates without trivializing the classifier (e.g., random predictions).
To understand why EO is not a desirable metric for fairness in target-group detection, consider a scenario with two marginalized demographic groups, Group-A and Group-B, that differ in how frequently they are targeted. When optimizing for EO, the model aligns true and false positive rates across both groups, ignoring the label imbalance. However, this alignment often requires redistributing errors, which disproportionately impacts the statistical minority group.
This is not ethically desirable for several reasons. First, “target group" refers to the group that a post is “directed at or about” [38], which is independent of the downstream task (e.g., whether the post is toxic). Being a frequent “target" does not inherently mean a group warrants prioritization. Second, a group that constitutes a small minority in society may be statistical minority in the data (across both toxic and non-toxic posts) simply by virtue of its size. This does not mean it should be de-prioritized.
Case Group Targeted Not Targeted TP FN FP TN Acc TPR FPR
I - Optimizing for Equalized Odds Group-A 100 100 80 20 30 70 0.75 0.80 0.30
Group-B 20 180 16 4 54 126 0.71 0.80 0.30
II - Optimizing for Accuracy Parity Group-A 100 100 77 23 23 77 0.77 0.77 0.23
Group-B 20 180 15 5 41 139 0.77 0.75 0.23
Table 2: Illustrative example of two marginalized groups with one (A) targeted more often than the other (B). Assume 200 total posts with Group-A having a higher base rate. To optimize for Equalized Odds (Case I), we set $TPR=0.80$ and $FPR=0.30$ across both groups. Alternatively, to optimize for Accuracy Parity (Case II), we set similar $Acc=0.77$ across both groups.
To illustrate how EO leads to a statistical minority groups being disadvantaged, we present a synthetic example (Table 2), where both groups have $TPR=0.8$ and $FPR=0.3$. Despite identical error rates, group-B exhibits lower accuracy $(0.71)$ due to its lower base rate and increased number of false positives. In contrast, optimizing for AP (Case II) equalizes overall accuracy across groups (0.77), reduces FPR for the group-B, and maintains similar TPR/FNR values. Note that the increase in FPR that group-B experiences under EO can lead to downstream harms. If the labels provided by such a target group detection module are used to train toxicity detection systems such as Gordon et al. [18], Fleisig et al. [15], Gupta et al. [20], then these models would likely fail at learning group-specific toxicity patterns affecting group-B, because the module often incorrectly flags this group as the target.
7 Experimental Setup
In this section, we describe the dataset used, the neural architecture, the baseline loss, and two baseline fairness losses used for comparing performance of the $GAP_{multi}$ loss. Implementation details regarding setup, data pre-processing, and related items can be found in Appendix B.
7.1 Dataset
We use the Measuring Hate Speech (MHS) corpus [38] as shared in the HuggingFace library [14]. The dataset has $135,556$ posts, sourced from YouTube, Twitter and Reddit, where each post is annotated for its target demographic group(s). We select posts targeting one or more demographic-race groups, irrespective of toxicity labels (which we do not use). Annotator labels indicate whether or not each post targets one or more of seven race groups: Asian, Black, Latinx, Middle-Eastern, Native-American, Pacific-Islander and White. Figure 1(a) shows how frequently each group is targeted. Figure 1(b) shows the number of groups targeted per post [^2] For groups of size six and seven, posts target a broader range of demographic categories, often using inclusive terms like ‘non-white’, as shown in Table 1. This explains the higher frequency of such posts compared to those targeting fewer than five groups. . We take a 80%-20% stratified train-test split of the data.
(a) Number of posts targeting each group.
(b) Number of posts targeting multiple groups.
Figure 1: Summary statistics of the MHS corpus [38] show the distribution of posts targeting demographic groups. The Black community is the statistical majority, while Native American and Pacific Islander are statistical minorities. Additionally, the dataset includes posts targeting multiple groups, reflecting its multi-group nature.
7.2 Neural Model and Baseline Measure
For our neural model (Figure 2), we use DistilBERT [39] to extract feature representations for posts [^3] One can replace the DistilBERT layer with any other feature representation without altering the rest of the model.. This is followed by layers of dense neuron connections with relu activation and added biases. The dense layers are of (512, 128, 64) neurons respectively, with a dropout rate of 0.1, to prevent model overfitting. The layers end in $G$ classification nodes, each with sigmoid activation, no biases and 0.5 classification threshold. Given the multi-label binary classification setup of our problem, we use weighted Binary Cross Entropy (wBCE) as our baseline loss to optimize Overall Error (OE) for the $G$ output nodes. Our experiments freeze the weights of the DistilBERT layer, hence the only trainable parameters are the nodes in the dense layers and the final classification layer.
Figure 2: Our multi target-group detection architecture. The model has shared parameters to learn both general and group-targeted language properties. The classification nodes each learn group-specific mappings.
7.3 Other Differentiable Fairness Measure
CLA. We compare to CLAss-wise equal opportunity (CLA) [42], a differentiable fairness loss. CLA seeks to balance False Negative Rate (FNR) across protected groups [9, 21] by minimizing error in absolute differences between error w.r.t. a label $(BCE(y))$ vs. error w.r.t. a label given the group $(BCE(y,g))$. We replace BCE with wBCE in Eq. 8 to compare fairly under label imbalance.
$CLA=BCE+\lambda\cdot\sum_{y\in C}\sum_{g\in G}|BCE(y,g)-BCE(y)|$ (8)
ADV. We also compare against a differentiable loss: ADV [50]. Since ADV only supports a 2-group setting, we present results with it in Appendix C.
7.4 Evaluation Measures Considered
For model evaluation, we consider balanced accuracy, hamming loss and macro-level measures to capture a comprehensive view of model performance for various groups across losses. These measures account for group imbalances, ensuring that the statistical minority groups are fairly evaluated, while macro-level measures provide an aggregate view that treats each group equally regardless of their size.
Balanced Accuracy (BA)
Balanced Accuracy provides an assessment of accuracy while accounting for label imbalance in datasets. It computes the average accuracy of each label, thereby offering a balanced perspective to account for the unequal label distribution. By considering both the sensitivity (TPR: true positive rate) and specificity (TNR: true negative rate) of each label, BA effectively captures the model’s ability to correctly classify instances across all labels, regardless of their prevalence.
$BA=(TPR+TNR)/2.0$ (9)
Average Balanced Accuracy (Avg. BA)
When optimizing accuracy across groups, we report the macro average over the group-specific BAs as a summary statistics. The Avg. BA (macro) treats each group equally, ensuring that the classifier’s performance is evaluated in a balanced manner across all demographic groups.
$Avg.\,BA=\frac{1}{G}\sum_{g=1}^{G}BA(g)$ (10)
Hamming Loss (HL)
This is a widely employed metric for assessing the performance of multi-label classifiers. Formally, for a dataset with $N$ instances and $G$ labels, the Hamming Loss (HL) quantifies the fraction of incorrectly predicted labels across all instances in the dataset, with $hamming(y_{i}(g),\hat{y}_{i}(g))$ as an indicator function of 1 if the $g$-th label for instance $i$ is incorrectly predicted and 0 otherwise. Specifically, it measures the average fraction of labels that are misclassified in comparison to the true label set. HL and Subset Accuracy Loss are comparable under small label cases [48], hence we just report HL.
$HL=\frac{1}{NG}\sum_{i=1}^{N}\sum_{g=1}^{G}hamming(y_{i}(g),\hat{y}_{i}(g))$ (11)
Other measures of interest
Although we are strictly optimizing for similar BA across groups, it is also important to assess the effect that this gain in fairness may have on other measures of interest, like Precision, Recall, F1. Given our multi-label setup, we report the macro variants (average over group-specific numbers) of the mentioned measures.
8 Results
Results show that $GAP_{multi}$ not only achieves better group-fairness by minimizing performance disparities impact across groups, but also maintains competitive overall accuracy. This highlights the effectiveness of $GAP_{multi}$ in addressing the unique challenges of multi-group target detection. All reported results in the main material are for the 7-group setting. For group performance on selected $\prescript{G\mkern-0.5mu}{}{C}_{2}$ isolated pairs, please refer to Appendix C.
8.1 Evaluation & Loss Performance
The values presented in Table 3 illustrate the achieved BA values across various groups over five runs for the test set, for the two baseline losses (OE and CLA) vs. $GAP_{multi}$.
Balanced Accuracy (BA) ($\uparrow$)
Loss Asian Black Latinx Middle Eastern Native American Pacific Islander White Max. Diff. ($\downarrow$) Avg. BA ($\uparrow$) #Best BA ($\uparrow$)
OE 80.31 86.91 81.07 84.87 64.99 67.91 75.01 21.9$\pm$1.3 77.29$\pm$0.29 2/7
CLA 82.51 85.34 81.67 84.62 73.91 74.92 80.44 11.4$\pm$0.8 80.49$\pm$0.14 0/7
$GAP_{multi}$ 83.18 83.86 83.47 83.42 78.95 78.32 82.58 5.5$\pm$0.5 81.97$\pm$0.13 5/7
Table 3: Balanced Accuracy (BA) achieved by each loss function (OE and CLA baselines vs. $GAP_{multi}$) over the 7 demographic groups (on test data) for five runs. For each loss, we also color which group exhibits the maximum and minimum BA values achieved for that loss over the 7 groups, with the difference between this maximum vs. minimum shown in the Max. Diff. column (for BA, we want to minimize this difference). $GAP_{multi}$ achieves a lower maximum difference (Max. Diff. = 5.54) than either baseline loss functions, evident from the visualization in Fig. 3. $GAP_{multi}$ also achieves the highest (macro) average BA across groups (Avg. BA = 81.97), and the best BA for most of the groups (#Best BA column = 5/7). We report the standard deviation of Max. Diff and Avg. BA over five runs, with CLA and $GAP_{multi}$ having similar deviation. $GAP_{multi}$ achieves the best fairness performance while retaining overall utility.
Recall that Overall Error (OE) corresponds to minimizing overall error without any fairness constraints, and CLA equalizes false negative rates. The Black group, which constitutes the statistical majority group in the dataset (as shown in Figure 1), exhibits the maximum BA values for the three losses, with relatively similar performance across all three. Meanwhile, performance for the other groups, and in particular for the Native American group, which is a statistical minority, varies starkly across the different optimization strategies. Moreover, the performance for the group that exhibits the minimum BA values for each loss shows that the worse-off group exhibits the best performance under $GAP_{multi}$.
Figure 3: Visualization of the BA values achieved by each loss over the 7 demographic groups. The maximum difference (Max. Diff.) between the maximum and minimum BA achieved for each loss across groups is also shown. See Table 3 for additional details. $GAP_{multi}$ performs best with lowest Max Diff. of 5.5, showcasing that it greatly reduces the disparities in performance across groups vs. other losses.
We show the maximum difference (Max. Diff.) between the groups in Table 3, which captures the gap in performance between worst vs. best performing groups. We see that solely optimizing for overall error (OE) fails to account for variations in group performance, resulting in the highest difference values (Max. Diff. of 21.9). In contrast, both the $GAP_{multi}$ and CLA approaches incorporate considerations of group performance alongside overall error, leading to substantially lower Max. Diff. compared to OE. Notably, $GAP_{multi}$ optimizes for balanced error rates across groups, exhibiting the smallest difference (Max. Diff. of 5.5).
We report the average BA (macro) score obtained in all three losses in Table 3, with $GAP_{multi}$ having the highest value (Avg. BA: 81.97). We hypothesize that by incorporating the group information in the loss during model training, both the group-informed losses (CLA and $GAP_{multi}$) are able to find better local optima compared to OE. The Best BA is a group measure indicating how many amongst the 7 groups are performing best across the three losses. For this measure, we observe OE to have the highest group BA values for 2 groups (Black, Middle Eastern), while $GAP_{multi}$ performs best for rest of the five groups. This also emphasizes the fact that $GAP_{multi}$ does not prioritize the performance of one group over others during optimization, thereby being the best performing loss for most of the groups.
Figure 4: Heatmap of pairwise absolute difference of BA across groups in test set as an indicator for bias and disparate impact. OE has the highest performance gap (Max Diff = 21.9) across groups as indicated by the extremes of color, not only across one group- pair but consistently across multiple group pairs. $GAP_{multi}$ has the least spread in pairwise error values (Max Diff = 5.5), evident from the flatness of color, indicating least disparate impact across groups.
[…]
The results for OE show that the (Black, Native American) pair has the largest Max. Diff. of 21.9. We also see other group pairs with wide performance disparity. Two key observations can be drawn. First, the Black group being the statistical majority group has a dominant performance gap over the Native American group, one of the statistical minorities. Second, even for group pairs that have similar prevalence in the data, e.g., the (Middle Eastern, White) pair, there can be performance disparities (BA gap of 9.9) because one group may simply be easier to identify than the other.
CLA, which optimizes for FNR, has improved performance over OE, however, since its optimization criteria (minimizing false negatives) does not align with the intended evaluation (having similar performance across groups), and its heatmap lies in-between that of OE and $GAP_{multi}$. In contrast, the $GAP_{multi}$ loss, which is explicitly designed to achieve similar (balanced) performance across groups while optimizing overall utility, shows substantially fewer extremes in group performance gaps. The heatmap reveals smoother transitions between groups, indicating equitable distribution of the classifier’s performance. Note that the extreme values of Maximum Difference (Max. Diff.) presented in Table 3 are reflected as outliers in these heatmaps, corresponding to specific group pairs.
OE CLA $GAP_{multi}$
Hamming Loss (HL) % ($\downarrow$) 7.65 12.10 6.85
Table 4: HL values across different losses for test data. Lower values are better. $GAP_{multi}$ optimizes for jointly over TPR and TNR, thereby achieving lowest values, indicating better classifier performance.
Hamming Loss (HL) (Eq. 11) values are shown in Table 4 as a summary statistic in our multi-label classification setup. The HL metric quantifies the average fraction of misclassified labels, with lower HL values indicating enhanced classifier performance. While the CLA loss (HL: 12.10) prioritizes minimizing FNR, which inherently involves asymmetry, there may be instances where it disproportionately optimizes for this aspect at the expense of other performance metrics. Consequently, CLA may exhibit poorer HL performance compared to OE loss (HL: 7.65). In contrast, the $GAP_{multi}$ loss maintains symmetry, resulting in a balanced trade-off while jointly minimizing TPR and TNR. As a result, $GAP_{multi}$ consistently achieves the lowest (HL: 6.85) values, indicative of superior classifier performance.
Loss $Prc_{macro}$ ($\uparrow$) $Rec_{macro}$ ($\uparrow$) $F1_{macro}$ ($\uparrow$)
OE 0.7083 0.5808 0.6383
CLA 0.5418 0.7143 0.6162
$GAP_{multi}$ 0.7854 0.6837 0.7310
Table 5: Summary statistics of other evaluation measures. Since CLA strictly optimizes for minimizing FNR, it indeed achieves the highest Recall (1 - FNR). However, this comes at the cost of losing out on Precision. Since $GAP_{multi}$ jointly minimizes both TPR and TNR, it performs best both in terms of Precision and F1 scores across three losses.
Table 5 shows summary statistics for Precision, Recall and F1 for multi-label classification setup at the macro level. The CLA approach, characterized by its emphasis on minimizing the False Negative Rate (FNR), inherently yields the highest Recall. However, this optimization strategy comes at the expense of Precision and F1, both being lower than OE. Our proposed $GAP_{multi}$ loss, designed to jointly minimize True Positive Rate (TPR) and True Negative Rate (TNR), emerges as the optimal performing loss function in terms of both Precision and F1 score across the evaluated losses.
8.2 Runtime Performance
Avg. Time
Per Epoch (s)
Epochs till
Convergence
Runtime
Total (s)
$\Delta$ (s)
w.r.t. OE
OE 154 21 3234 0
CLA 158 41 6478 3244
$GAP_{multi}$ 163 27 4401 1167
Table 6: Loss Runtime Analysis. While OE takes the least time, $GAP_{multi}$ gains more in optimizing performance across groups for an extra of $\sim 9s$ per epoch. The smoothness of $GAP_{multi}$ loss (27 epoch) also allows faster convergence compared to CLA (41 epoch).
Table 6 shows runtime performance results over the training dataset of 36k posts, where we report the average time per epoch, number of epochs till convergence, total runtime, and the extra time ($\Delta$) for losses compared to OE.
Since OE optimizes for weighted Binary Cross Entropy (wBCE), it takes the least time per epoch and the least number of epochs to converge. $GAP_{multi}$ takes all the steps for computing the overall loss in addition to calculating $\prescript{G\mkern-0.5mu}{}{C}_{2}$ pairwise losses and the balanced loss. Hence, $GAP_{multi}$ takes additional compute time for solving its intended optimization. The same argument holds true for CLA, since all of them are regularized single objective variants. Although $GAP_{multi}$ does the $\prescript{G\mkern-0.5mu}{}{C}_{2}$ extra computation, hence the extra runtime ($\Delta$: 1167s), it is not that significant (extra 9s per epoch) w.r.t. to OE, while gaining much more in terms of optimization improvement, because the pairwise error computations are all done in parallel. While OE and $GAP_{multi}$ operate on smooth losses their convergence epoch is relatively fast ($\sim$21, $\sim$27). CLA uses a 1-norm loss (Eq. 8), hence the empirical loss surface is not as smooth as the previous two. As such, it is observable that CLA on average takes more epochs ($\sim$41) to converge with a higher $\Delta$.
9 Discussion
Multi-group target detection brings various challenges and complexities, particularly in balancing the trade-offs between utility and group fairness, as our findings highlight. Achieving fairness in this setting involves mitigating disparities across diverse demographic groups, which is further complicated by various issues: data imbalance, base rate disparities, and the multi-label nature of the task. In this section we delve into some broader implications of our model formulation and empirical results. Additionally, we discuss its applicability in other fairness-sensitive domains and models, and the opportunities it presents for future work.
Fairness measures must be carefully selected to align with the requirements of the task, context, and stakeholders involved. In target-group detection, where posts may simultaneously target multiple groups, the symmetric error costs motivates a fairness measure that enforces equitable treatment across all groups. In this context, we adopted Accuracy Parity (AP) as the fairness criterion, as it directly addresses disparities in group-level predictive accuracy. AP measures the extent to which no group, particularly those with smaller representation or lower base rates, experiences disproportionate harm. This is accomplished using the $GAP_{multi}$ loss function, which equivalently minimizes for AP. The formulation enforces that fairness is incorporated directly into the model training process, enabling balanced performance without requiring additional pre- or post-processing steps.
Ensuring fairness across a growing number of demographic groups $G$ poses computational challenges, particularly for methods relying on sequential global metrics like deviations from mean. In contrast, our $GAP_{multi}$ formulation leverages a pairwise structure, computing error deviations between all group pairs $\prescript{G\mkern-0.5mu}{}{C}_{2}$ independently. This design enables parallelization, integrating seamlessly with modern deep learning pipelines and exploiting GPU acceleration for efficiency. Despite the quadratic growth in terms of $G$, the $GAP_{multi}$ loss retains its convergence properties, ensuring stable optimization even for high group cardinalities, making it robust for real-world multi-group scenarios.
For the 7-group setting in Table 3, we see that by optimizing for group-related errors alongside overall error, both $GAP_{multi}$ and CLA driven classifiers achieve better overall performance (in terms of Avg. BA), and also achieve their intended group-specific objective. Given our multi-label classification where the groups serve as the labels, we hypothesize that the group indicator in each classification node gives an extra feature dimension. This likely enables the model to learn something more about the data than it could without the group-specific information. By incorporating group-associated terms, both $GAP_{multi}$ and CLA have a modified surface compared to OE, allowing convergence to a better optima. We see this pattern emerging in some of the 2-group setting: for example, the Latinx vs. Middle Eastern group-pair in Appendix C.
Model Multiplicity [4] refers to the phenomenon where different models achieve similar overall utility but produce varying predictions. With target-group detection, classifiers trained using CLA and $GAP_{multi}$ exhibit similar overall accuracy but diverge significantly in their group-level performance. In this scenario, AP as a fairness measure provides valuable insight by highlighting the amount of bias across groups, with $GAP_{multi}$ having the least pairwise performance gap across groups.
Incorporating symmetric error considerations, where type-I and type-II errors are considered equally harmful, not only enhances the fairness of our multi target-group detection model but also opens new avenues to explore similar scenarios and challenges in other domains.
For example, this type of error is associated with tasks of inferring preferences often arising in recommender systems such as Netflix or Spotify. Such systems infer user preferences to suggest movies or songs, which may significantly vary across different demographic groups. A symmetric error approach suggests that recommendations are not skewed toward the preferences of the statistical majority group. By recognizing and addressing the symmetric nature of error costs across different groups with similar intent, we can achieve a balanced perspective on fairness and more equitable outcomes in algorithmic decision-making processes and avoid any portability traps [41].
Fair target-group detection is also relevant in the context of training and deploying Large Language Models (LLMs), such as role prompting LLMs to respond as as a specific character or persona (essentially guiding its tone, style, and reasoning approach) [46, 37, 43]. Studies have found that LLMs contain bias with respect to protected characteristics such as gender and race [34, 27]. Integrating explicit target group detection into LLM training pipelines, particularly through reinforcement learning from human feedback (RLHF), could mitigate bias by dynamically aligning model outputs with fairness constraints across demographic groups.
10 Conclusion
We have proposed a fairness-aware formulation of multi target-group detection, focusing on realistic scenarios where social media posts may target multiple demographic groups. Motivated by the symmetric harm of misclassification in this setting, we adopted Accuracy parity (AP) as a fairness criteria and proposed $GAP_{multi}$, a loss function that minimizes bias in a multi-group setting. Our method jointly optimizes for group fairness and utility, aligning training dynamics with evaluation measures. Through theoretical analysis, we demonstrated the incompatibility between Accuracy Parity and Equalized Odds, revealing trade-offs that are important to consider in group identity sensitive NLP tasks. Empirical evaluations on the MHS corpus, a large multi-label dataset with fine-grained demographic annotations, show that our approach consistently reduces bias across groups while retaining overall utility. By providing the first evaluation for this task, our work advances model development in fairness-aware target-group detection, contributing towards the development of safer and inclusive online spaces.
Appendix A Impossibility Proofs
Assume two groups $A$ and $B$. Let the number of positive and negative examples in group $A$ be $P_{A}$ and $N_{A}$ respectively. Similarly, $P_{B}$ and $N_{B}$ for group $B$ respectively. Proportion of positive to negative examples in a group, i.e., $\delta_{A}=\frac{P_{A}}{N_{A}}$ and $\delta_{B}=\frac{P_{B}}{N_{B}}$.
Since we are satisfying Equalized Odds, we have:
$TPR_{A}=TPR_{B}=TPR$
$FPR_{B}=FPR_{B}=FPR$
Proof of Theorem 1.
From basic definitions, we know the following:
True Positive Rate (TPR) $=\frac{TP}{TP+FN}$
False Positive Rate (FPR) $=\frac{FP}{FP+TN}$
Accuracy (Acc) $=\frac{TP+TN}{TP+TN+FP+FN}$
$=\frac{TP+TN}{P+N}$
For Group A, we have the following:
$TP_{A}$ $=TPR*P_{A}$
$FN_{A}$ $=P_{A}-TP_{A}=P_{A}*(1-TPR)$
$FP_{A}$ $=FPR*N_{A}$
$TN_{A}$ $=N_{A}-FP_{A}=N_{A}*(1-FPR)$
$\therefore Acc_{A}$ $=\frac{TP_{A}+TN_{A}}{P_{A}+N_{A}}$
$=\frac{TPR*P_{A}+N_{A}*(1-FPR)}{P_{A}+N_{A}}$ (12)
For Group B, we have the following:
$TP_{B}$ $=TPR*P_{B}$
$FN_{B}$ $=P_{B}-TP_{B}=P_{B}*(1-TPR)$
$FP_{B}$ $=FPR*N_{B}$
$TN_{B}$ $=N_{B}-FP_{B}=N_{B}*(1-FPR)$
$\therefore Acc_{B}$ $=\frac{TP_{B}+TN_{B}}{P_{B}+N_{B}}$
$=\frac{TPR*P_{B}+N_{B}*(1-FPR)}{P_{B}+N_{B}}$ (13)
Remark.
A fairness problem aiming to simultaneously satisfy Equalized Odds and Accuracy is only feasible when the base rates are equal across groups.
This remark extends Kleinberg et al. [26], showing that overall accuracy and EO (group-fairness) are at odds with each other for most practical cases when base rates differ across demographic groups ($\delta^{*}_{A}\neq\delta^{*}_{B}$), s.t. one needs to be sacrificed for the gain of the other. Enforcing EO may thus come at the cost of significant accuracy degradation, particularly in imbalanced datasets. Hence, we can observe that group accuracies are directly dependent on the base rates of each group. Therefore, enforcing EO on that would cause one or both groups to deviate from their optimal accuracy levels. For example if $\delta_{A}>\delta_{B}$, then to equalize TPR, group B must increase its TPR disproportionately, which decrease its accuracy levels.
Expanding on Eq. 12 and Eq. 13, we have the following two scenarios:
Special Case 1: Base rates across groups are equal, i.e., $\frac{P_{A}}{N_{A}}=\frac{P_{B}}{N_{B}}$, $P_{A}=\alpha P_{B},N_{A}=\alpha N_{B}$, $\forall\alpha\in(0,\infty)$.
Under this special setting of equalized base rates, Eq. 12 and Eq. 13 becomes the same:
$Acc_{A}$ $=\frac{TPR*P_{A}+N_{A}*(1-FPR)}{P_{A}+N_{A}}$
$=\frac{TPR*\alpha P_{B}+\alpha N_{B}*(1-FPR)}{\alpha P_{B}+\alpha N_{B}}$
$=\frac{\alpha(TPR*P_{B}+N_{B}*(1-FPR))}{\alpha(P_{B}+N_{B})}$
$=Acc_{B}$
Special Case 2: Base rates are not equal, but $TPR$ and $FPR$ sum to one. $TPR+FPR=1$. Under this special setting of unequalized base rates, Eq. 12 and Eq. 13 becomes the same. This is equivalent to random predictions by the classification model.
$Acc_{A}$ $=\frac{TPR*P_{A}+N_{A}*(1-FPR)}{P_{A}+N_{A}}$
$=\frac{TPR*P_{A}+N_{A}*TPR}{P_{A}+N_{A}}$
$=\frac{TPR(P_{A}+N_{A})}{P_{A}+N_{A}}$
$=TPR$
$Acc_{B}$ $=\frac{TPR*P_{B}+N_{B}*(1-FPR)}{P_{B}+N_{B}}$
$=\frac{TPR*P_{B}+N_{B}*TPR}{P_{B}+N_{B}}$
$=\frac{TPR(P_{B}+N_{B})}{P_{B}+N_{B}}$
$=TPR$
Case 1 demonstrates that AP and EO are mathematically compatible only when base rates are equal across groups, an assumption that rarely holds in practice, as most real-world datasets exhibit inherent demographic imbalances. Case 2 reveals a more troubling scenario: even with unequal base rates, simultaneous satisfaction of AP and EO forces the classifier into degenerate behavior, reducing it to random guessing (i.e., predicting uniformly at chance level). This shows a fundamental tension, i.e., strictly enforcing both criteria under realistic (imbalanced) conditions either violates feasibility (Case 1) or renders the model useless (Case 2).
Proof of Theorem 2. The False Positive and Negative Error Disparity (FPNED) [7] can be broken down into FPED and FNED.
$FPED=\sum_{G}|FPR-FPR_{g}|$
$FNED=\sum_{G}|FNR-FNR_{g}|$
Only when $FPR_{A}=FPR_{B}$ and $FNR_{A}=FNR_{B}$, does $FPED=0$ and $FNED=0$. Eq. 12 and Eq. 13 can be rephrased as:
$Acc_{A}$ $=\frac{TP_{A}+TN_{A}}{P_{A}+N_{A}}$
$=\frac{(1-FNR)*P_{A}+(1-FPR)*N_{A}}{P_{A}+N_{A}}$
$Acc_{B}$ $=\frac{TP_{B}+TN_{B}}{P_{B}+N_{B}}$
$=\frac{(1-FNR)*P_{B}+(1-FPR)*N_{B}}{P_{B}+N_{B}}$
For the accuracies to be same, we need $Acc_{A}=Acc_{B}$. Equating and rearranging the terms we have:
$P_{A}*(P_{B}+N_{B})+N_{A}*(P_{B}+N_{B})$
$=P_{B}*(P_{A}+N_{A})+N_{B}*(P_{A}+N_{A})$
which only holds when $\frac{P_{A}}{N_{A}}=\frac{P_{B}}{N_{B}}$, showing that AP and FPNED are only equivalent when base rates are equal.
Appendix B Experimental Details
B.1 Setup
Experiments use a Nvidia 2060 RTX Super 8GB GPU, Intel Core i7-9700F 3.0GHz 8-core CPU and 16GB DDR4 memory. We use the Keras [8] library on a Tensorflow 2.0 backend with Python 3.7 to train the networks in this paper. For optimization, we use AdaMax [25] with parameters (lr=0.001) and $1000$ steps per epoch. We also setup an EarlyStopping criteria in case the with a min_delta of 1e-4 and patience of 5.
B.2 Data Pre-processing
From the set of all posts, we filter out ones that target at least one or more demographic group(s), resulting in a dataset size of 44,816 posts. From there we perform a 80%-20% split for training vs. testing data, for all cluster sizes that allowed for a SkLearn’s [36] stratified sampling following the label split ratio. This resulted in a train set of size 36,412 posts and test set of size 8,404 posts.
B.3 Loss
Another fairness loss [50] is an adversarial approach to demoting unfairness, which we denote as ADV. It seeks to provide false positive rate (FPR) balance [9] across groups, otherwise known as predictive equality (ibid.)
$ADV=\beta\cdot BCE+(1-\beta)\cdot(adversary(y,g)-0.5)$ (14)
Appendix C Pairwise Group Classifiers
We present the performance across four losses (OE, CLA, ADV and $GAP_{multi}$) for some of the 2-group case to show different scenarios and performances. Best values are bolded (higher for BA, lower for Diff.). Since ADV cannot accommodate multiple groups, we were unable to report ADV numbers for the 7-group case in the main material. ADV [50] is an approximate adversarial loss for balancing FPR rates across groups. We notice similar issues of convergence instability as they observed in Xia et al. [50] as well. Consequently, let ADV run for a fixed epochs and report the best BA value achieved over iterations.
Balanced Accuracy (BA)
Loss Latinx Middle Eastern Avg. BA Diff.
OE 90.98 83.67 87.33 7.31
CLA 91.52 88.73 90.12 2.79
ADV 91.04 84.20 87.62 6.84
GAP 92.34 91.79 92.06 0.55
Table 7: Optimizing across groups results in improving overall error, indicating that the group label provides an extra dimension for the loss to stabilize at a better local optima.
Balanced Accuracy (BA)
Loss Asian Black Avg. BA Diff.
OE 86.59 92.32 89.46 5.73
CLA 87.02 92.22 89.62 5.20
ADV 87.49 92.10 89.79 4.61
GAP 89.65 90.82 90.23 1.17
Table 8: Performance varies across groups due to their population size, where the statistically major group dominates.
Balanced Accuracy (BA)
Loss White Latinx Avg. BA Diff.
OE 88.89 82.19 85.54 6.70
CLA 88.20 85.55 86.87 2.65
ADV 88.18 82.19 85.18 5.99
GAP 88.36 86.23 87.29 2.13
Table 9: Groups have similar population, but performance varies due to one group being more difficult to model.